# GlycanML: A Multi-Task and Multi-Structure Benchmark for Glycan Machine Learning

### v2

@@ 5.3 Benchmark Results on Multi-Task Learning
  • MTL is more helpful for difficult tasks on RGCN. From domain-level to species-level classification (i.e., from 4-way classification to 1,737-way classification), the task difficulty monotonically increases. Under RGCN, we can observe more benefits of MTL on difficult tasks. For example, the DTP method outperforms single-task learning on order-, family, genus- and species-level classification (i.e., tasks with more categories), while the DTP method is inferior on kingdom-, phylum- and class-level classification (i.e., tasks with fewer categories). Therefore, in MTL with RGCN, more attention is paid to difficult tasks, leading to better performance on these tasks.
@@
5.4 The Effect of Multi-Task Learning on Glycan Function Prediction
@@ 6 Conclusions and Future Work
In this work, we build a comprehensive benchmark GlycanML for glycan machine learning. It consists of diverse types of glycan understanding tasks, including glycan taxonomy prediction, glycan immunogenicity prediction, glycosylation type prediction, and protein-glycan interaction prediction. In GlycanML, we support two representation methods of glycans, i.e., glycan tokenized sequences and glycan planar graphs. Additionally, on eight highly correlated glycan taxonomy prediction tasks, we set up a testbed GlycanML-MTL to compare different multi-task learning (MTL) algorithms. Also, we study how taxonomy prediction can boost other three function prediction tasks by MTL. According to the benchmark results, multi-relational graph neural networks (GNNs) show great promise for glycan modeling, and well-designed MTL methods can further boost model performance.
Limitations and future work. The current GlycanML benchmark is limited to modeling glycan sequences and 2D glycan graphs, without benchmark datasets of 3D glycan structures. Therefore, in the future, we will go beyond sequence- and 2D-graph-based datasets and methods to 3D-structure-based ones [5, 16]. Also, we will work along with the community to further promote the efforts of MTL for better glycan understanding.
7 Broader Societal Impacts
This work aims at building a comprehensive benchmark for glycan machine learning, incorporating diverse types of glycan understanding tasks, multiple representation methods of glycan structures, and a testbed for multi-task learning algorithms. By evaluating on the proposed benchmark, we can judge the general effectiveness of a machine learning model on predicting different glycan functions. Therefore, such a benchmark can promote the application of machine learning methods to various real-world glycan-related tasks, such as vaccine design [29] and cancer research [52].
However, we should not ignore the potential harmful aspects brought by glycan machine learning models developed on our benchmark, e.g., designing vaccines with severe adverse reactions. To mitigate such risks, our future works will encourage the responsible usage of the GlycanML benchmark for real-world problems.

### v3

@@ 5.3 Benchmark Results on Multi-Task Learning
Method Domain Kingdom Phylum Class Order Family Genus Species Mean Macro-F1 Backbone encoder: Shallow CNN Single-Task 0.629(0.005) 0.559(0.024) 0.388(0.024) 0.342(0.020) 0.238(0.016) 0.200(0.014) 0.149(0.009) 0.115(0.008) 0.327(0.003) N-MTL 0.619(0.009) 0.549(0.040) 0.333(0.029) 0.344(0.015) 0.213(0.015) 0.167(0.014) 0.150(0.008) 0.121(0.010) 0.312(0.010) GN 0.646(0.043) 0.497(0.045) 0.344(0.064) 0.312(0.002) 0.233(0.042) 0.206(0.027) 0.175(0.038) 0.137(0.038) 0.319(0.012) TS 0.642(0.031) 0.609(0.059) 0.393(0.061) 0.355(0.013) 0.249(0.019) 0.187(0.009) 0.141(0.010) 0.113(0.009) 0.336(0.014) UW 0.632(0.005) 0.620(0.046) 0.409(0.025) 0.320(0.011) 0.230(0.012) 0.179(0.012) 0.152(0.001) 0.118(0.008) 0.332(0.005) DWA 0.619(0.012) 0.536(0.018) 0.354(0.013) 0.330(0.041) 0.218(0.011) 0.152(0.018) 0.138(0.015) 0.099(0.012) 0.305(0.013) DTP 0.633(0.004) 0.556(0.023) 0.341(0.041) 0.303(0.028) 0.213(0.034) 0.171(0.010) 0.145(0.011) 0.107(0.006) 0.309(0.017) Backbone encoder: RGCN Single-Task 0.633(0.001) 0.647(0.054) 0.462(0.033) 0.373(0.036) 0.251(0.012) 0.203(0.008) 0.164(0.003) 0.146(0.004) 0.360(0.009) N-MTL 0.641(0.001) 0.646(0.066) 0.402(0.042) 0.376(0.054) 0.231(0.025) 0.190(0.007) 0.154(0.018) 0.119(0.007) 0.345(0.020) GN 0.626(0.002) 0.612(0.050) 0.379(0.011) 0.364(0.017) 0.233(0.010) 0.169(0.006) 0.148(0.009) 0.119(0.006) 0.331(0.005) TS 0.640(0.003) 0.652(0.050) 0.461(0.013) 0.388(0.031) 0.267(0.024) 0.209(0.010) 0.174(0.002) 0.151(0.002) 0.368(0.016) UW 0.632(0.014) 0.568(0.045) 0.395(0.042) 0.386(0.045) 0.245(0.015) 0.199(0.009) 0.165(0.005) 0.133(0.012) 0.341(0.019) DWA 0.587(0.095) 0.616(0.112) 0.400(0.035) 0.379(0.024) 0.230(0.017) 0.177(0.023) 0.151(0.015) 0.110(0.005) 0.331(0.032) DTP 0.647(0.019) 0.637(0.083) 0.399(0.069) 0.370(0.023) 0.266(0.031) 0.228(0.039) 0.190(0.043) 0.158(0.035) 0.362(0.007)
5.4 Effect of Multi-Task Learning on Glycan Function Prediction
@@ 6 Conclusions and Future Work
In this work, we build a comprehensive benchmark GlycanML for glycan machine learning. It consists of diverse types of glycan understanding tasks, including glycan taxonomy prediction, glycan immunogenicity prediction, glycosylation type prediction, and protein-glycan interaction prediction. In GlycanML, we support two representation methods of glycans, i.e., glycan tokenized sequences and glycan planar graphs, enabling glycan modeling with sequence encoders and graph neural networks (GNNs). Additionally, on eight highly correlated glycan taxonomy prediction tasks, we set up a testbed GlycanML-MTL to compare different multi-task learning (MTL) algorithms. Also, we study how taxonomy prediction can boost other three function prediction tasks by MTL. According to the benchmark results, multi-relational GNNs show great promise for glycan modeling, and well-designed MTL methods can further boost model performance.
In the future, we will apply glycan machine learning models to important real-world glycan-related tasks. For example, for vaccine design, we can employ well-performing immunogenicity predictors and protein-glycan interaction predictors to virtually screen glycan candidates that can most effectively induce immune response.
Appendix A Details of Model Architectures
Model Input Layer Hidden Layers Output Layer #Params. Sequence encoders Shallow CNN 128-dim. token embedding 2 $\times$ 1D conv. layers (hidden dim.: 128; max pooling over all tokens 191.7K kernel size: 5; stride: 1; padding: 2) ResNet 512-dim. token embedding 8 $\times$ residual blocks (hidden dim.: 512; attentive weighted sum over all tokens 11.4M + 512-dim. positional embedding kernel size: 3; stride: 1; padding: 1) LSTM 640-dim. token embedding 3 $\times$ bidirectional LSTM layers weighted sum over all tokens 26.7M (hidden dim.: 640) + linear (output dim.: 640) + Tanh Transformer 512-dim. token embedding 4 $\times$ Transformer blocks (hidden dim.: 512; linear (output dim.: 512) + Tanh 21.4M + 512-dim. positional embedding #attn. heads: 8; activation: GELU) upon [CLS] token Homogeneous GNNs GCN 128-dim. node embedding 3 $\times$ GCN layers concat. mean & max pooling 67.8K GAT 128-dim. node embedding 3 $\times$ GAT layers (#attn. heads: 2) concat. mean & max pooling 69.4K GIN 128-dim. node embedding 3 $\times$ GIN layers concat. mean & max pooling 117.4K Heterogeneous GNNs MPNN 128-dim. node & edge embedding 3 $\times$ MPNN layers Set2Set pooling (#steps: 3) 4.0M RGCN 128-dim. node embedding 3 $\times$ RGCN layers concat. mean & max pooling 4.2M CompGCN 128-dim. node embedding 3 $\times$ CompGCN layers concat. mean & max pooling 150.4K All-Atom Molecular Encoders Graphormer 768-dim. node embedding 8 $\times$ Graphormer layers concat. mean & max pooling 28.5M GraphGPS 512-dim. node & edge embedding 6 $\times$ GraphGPS layers concat. mean & max pooling 17.4M
We provide the detailed architectures of baseline models in Table 5. For the input layer of each model, we employ an embedding layer to get the representation of each token/node; an additional positional embedding layer is used by ResNet and Transformer to represent sequential information; MPNN adopts an additional edge embedding layer to represent glycosidic bonds. We follow the implementation of different sequence and graph encoding layers in TorchDrug [68] to construct hidden layers. For the output layer of sequence encoders, we follow the implementation in TorchDrug to readout a glycan-level representation from token-level representations. We use the concatenation of mean and max pooling as the output layer of GCN, GAT, GIN, RGCN, CompGCN, Graphormer and GraphGPS, for its superior performance in our experiments. Following the original design of MPNN, a Set2Set pooling [56] serves as its output layer. We follow the default architecture of Graphormer and GraphGPS in their original work.